OpenStudy (anonymous):

write an equation of a line in slope intercept form going through the points (- 2, 0) and (4, - 3). Show your work

OpenStudy (harsimran_hs4):

given a line passes through (x1, y1) and (x2, y2) then slope(m) = y2-y1/(x2-x1) and eq of line is y-y1 = m (x - x1) or y-y2 = m(x - x2)

OpenStudy (anonymous):

i dont know how to do any of this v~v

OpenStudy (harsimran_hs4):

v~v ?

OpenStudy (anonymous):

its a face

OpenStudy (harsimran_hs4):

then what have you studied till now about lines ?

OpenStudy (anonymous):

first time.. no idea..

OpenStudy (harsimran_hs4):

then maybe you should refer to a standard text and get to know various type of equation of lines

OpenStudy (anonymous):

idk wat your talkn about

OpenStudy (harsimran_hs4):

i mean first go and read some material about lines and then you will have better understanding and you will be able to solve this problem

OpenStudy (anonymous):

i only have 12 days... i dont have time...

OpenStudy (harsimran_hs4):

you have it will only take about 1-2 hrs

OpenStudy (anonymous):

im only here till 1:00 and i have other problems to do as well

OpenStudy (harsimran_hs4):

then read my 1st comment/post it`s very clear if you follow it step by step

OpenStudy (harsimran_hs4):

ok then read my first comment and tell me the value of m you get ?

OpenStudy (anonymous):

thts just it i dont know how... no matter how clear it shows

OpenStudy (harsimran_hs4):

(-2, 0) = (x1, y1) (4, -3) = (x2, y2) now tell me x1- x2 and y1- y2

OpenStudy (anonymous):

ik at x's and the y's are i just dont know how to work it out

OpenStudy (harsimran_hs4):

(-2, 0) = x1, y1 means x1 is -2 and y1 = 0

OpenStudy (anonymous):

yes ik tht part i jus dont know how to work it out

OpenStudy (marissalovescats):

Well first off because you have 2 points, you need to find your slope. This will be your m. y2-y1/x2-x1 AND then you only have to use one set of ordered pairs to plug into the equation of y-y1=m(x-x1) Whatever ordered pair you choose will be your x1 and y1. Let me know if you need me to show you how to do all of this. I would be more than happy to teach you.

OpenStudy (anonymous):

First we need to find the slope(m) m = (y2 - y1( / (x2 - x1) y2 = -3, y1 = 0 x2 = 4, x1 = -2 now lets sub our info into the formula m = (-3 - 0) / (4 - (-2) can you finish ?

OpenStudy (anonymous):

whats -3 - 0 and what 4 -(-2) ???

OpenStudy (anonymous):

-3 and 6

OpenStudy (anonymous):

the formula would make it -3/6 which equals -1/2 -1/2 is the slope. Now we will use y - y1 = m(x - x1) m = -1/2 as for x1 and y1 , you can pick either of your points....I will pick (-2,0) so x1 = -2and y1 = 0 now sub your info in the formula.. y - y1 = m(x - x1) y - 0 = -1/2(x - (-2) can you do this ? First distribute the -1/2 through the parenthesis.

OpenStudy (anonymous):

Let me rephrase that...do what needs to be done inside the parenthesis, then distribute through the parenthesis.

OpenStudy (anonymous):

you haven't gotten to the answer yet.....

OpenStudy (anonymous):

sorry i was readn it n tryn to understand it lol sowwy

OpenStudy (anonymous):

y - 0 = -1/2(x - (-2) can you do this ?

OpenStudy (anonymous):

if you told me how lol

OpenStudy (anonymous):

look inside the parenthesis.....(x -(-2)....what does -(-2) equal ?

OpenStudy (anonymous):

im not the sharpest knife in the crayon box....:/

OpenStudy (anonymous):

y - 0 = -1/2(x - (-2) y = 0 - 1/2(x + 2) now distribute y = 0 - 1/2x - 1 are you with me so far ?

OpenStudy (anonymous):

oops....I should be putting y - 0 not y = 0...sorry about that

OpenStudy (anonymous):

@MAYSTEPH .....lets finish this

OpenStudy (anonymous):

Let me know when your ready

OpenStudy (anonymous):

ok idk

OpenStudy (marissalovescats):

The equation is right in front of you. You can do it.

OpenStudy (anonymous):

how :(

OpenStudy (anonymous):

ok...I will finish because I have to go y - 0 = -1/2x - 1 y = -1/2x - 1 <--your equation

OpenStudy (anonymous):

do you have any questions ?

OpenStudy (anonymous):

not if you gotta go lol

OpenStudy (anonymous):

I have a few minutes.....do you have any questions ?

OpenStudy (anonymous):

You really need to learn this...it is not really that hard if you put a little effort into it...you do not have to be a math whiz to do this

OpenStudy (anonymous):

lol aloooot but idk how to ask lol

OpenStudy (anonymous):

I can solve these problems but I am not good at graphing...so if they are graphing questions, you will need to ask someone else

OpenStudy (anonymous):

just get the formula's and sub in your info....and solve. I have to go now....I wish you the best of luck :) Have a nice day

OpenStudy (anonymous):

thankyouuuuu